OpenStudy (anonymous):

Simplify the exponential expression. 3x^10 / x^6

OpenStudy (anonymous):

I'm not sure how to simplify this...\[\frac{ 3x ^{10} }{ x^6 }\]

OpenStudy (anonymous):

\(\Large \frac{3x^{10} }{ x^6}=3\times \frac{x^{10}}{x^6}=3\times x^{10-6}=3x^4\)

OpenStudy (igreen):

Divide \(3 \div 1\) SUBTRACT the exponents. What's your answer? @AngelCriner

OpenStudy (igreen):

@JungHyunRan You're not supposed to give direct answers.

OpenStudy (anonymous):

Thank you both, and it's okay. I solved it the same way, but I wasn't sure if I was correct. I was though yay and thank you both

OpenStudy (dawnr):

why not? o.O

OpenStudy (anonymous):

because understanding the answer is better than getting the answer o.o and people will just cheat if they get the answer rather than work it out I suppose

OpenStudy (igreen):

^^

OpenStudy (igreen):

It's part of their "code-of-conduct". I learned that the hard way.. :l

OpenStudy (dawnr):

well she explained the whole thing ..a person would need to be a retard not to understand it xD

OpenStudy (dawnr):

the sites or?

OpenStudy (igreen):

Yes, but they still want the person who asked the question to FIND OUT THE ANSWER FOR THEM SELF.

OpenStudy (igreen):

I think it's dumb too..

OpenStudy (dawnr):

oh didn't know that..that explains things xDDD

OpenStudy (igreen):

'Cause sometimes it takes 10 minutes to get 1 medal because the person can't understand what you're saying..

OpenStudy (dawnr):

yes i know..i usually solve the whole thing when i'm explaining something..didn't know of this xD

OpenStudy (igreen):

Lol

OpenStudy (igreen):

I did that before and got banned..

OpenStudy (dawnr):

wow ..thanks for telling me! :D

OpenStudy (dawnr):

can you help me solve something then?! :D

OpenStudy (igreen):

Solve what? @DawnR